OpenStudy (anonymous):

Find the slope-intercept form for the line satisfying the following conditions. Passing through (3,-2) and (2,-1)

OpenStudy (mathstudent55):

The slope-intercept form is y = mx + b, where m = slope, and b = the y-intercept, or point (0, b) First you need to find the slope: The slope is the rise over run, or difference in y-values divided by the difference in x-values. For the given points, the y-values are -2 and -1. The x-values are 3 and 2. The slope = m = (-2 - (-1))/(3 - 2) = (-2 + 1)/1 = -1/1 = -1 You equation will look like y = -x + b To find b, one method is to insert the x- and y-coordinates of one of the given points in for x and y, and solve for b. Let's use point (3, -2) y = -x + b -2 = -(3) + b -2 = -3 + b b = 1 The equation is: y = -x + 1